OpenStudy (anonymous):

Can someone please help me?? Haley has deposited $816 in a savings account that earns interest at a rate of 2.9% compounded daily. What will the account balance be in nine years? $1,055.43 $2,366.40 $839.66 $1,059.34

jimthompson5910 (jim_thompson5910):

A = P(1 + r/n)^(n*t) A = 816(1 + 0.029/365)^(365*9) A = ???

OpenStudy (anonymous):

1059.3428? So D?

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

Thank you so much <333

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

could you help me with another, please?

jimthompson5910 (jim_thompson5910):

ok, one more

OpenStudy (anonymous):

Thank you! I appreciate it so much. <3 Fabian has a savings account balance of $1,374.73. The interest rate on the account is 3.9% compounded twice a year. If he opened the account nine years ago, what was the value of his initial deposit? $1,154.43 $1,322.23 $352.49 $971.06

jimthompson5910 (jim_thompson5910):

ok one sec

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

A = P(1 + r/n)^(n*t) 1374.73 = P(1 + 0.039/2)^(2*9) now you want to solve for P to get your answer does that help?

OpenStudy (anonymous):

Yes it does! thank you! So the answer is D $971.06 ? :)

jimthompson5910 (jim_thompson5910):

very good, you got the hang of this

OpenStudy (anonymous):

You really helped me understand it. I was confused about the formulas but now I'm not. Thanks again, you were a great help. I got an A on my test. :)